# On manifolds with almost non-negative Ricci curvature and integrally-positive $k^{th}$-scalar curvature

### v2

@@ 4 Thin metric spaces
$\quad\phi(t^{b})\text{ has a projection on }b\text{ lying in }I^{b}\text{ for every }b\in\{\gamma_{1},\gamma_{2},l\}\Big{\}}.$
Notice that $A\neq\emptyset$ since $\gamma_{1}+\gamma_{2}+l\in A$. We claim that if $\{\phi_{i}\}_{i\in\mathbb{N}}\subset A$ and $\phi_{i}\to\phi_{\infty}$ uniformly, then $\phi_{\infty}\in A$. First of all $\phi_{\infty}$ is $1$-Lipschitz and satisfies $\phi_{\infty}(0)=\phi_{\infty}(\gamma_{1}+\gamma_{2}+l)$ trivially. Let $b\in\{\gamma_{1},\gamma_{2},l\}$. Consider for every $i\in\mathbb{N}$ the corresponding $t_{i}^{b}$. Modulo passing to a subsequence, there exists $t_{\infty}^{b}\in[0,\mathrm{L}(\gamma_{1}+\gamma_{2}+l)]$ such that $t_{i}^{b}\to t_{\infty}^{b}$. By Lemma 4.4, $\phi_{\infty}(t_{\infty}^{b})$ has a projection on $b$ lying in $I^{b}$. It follows that $\phi_{\infty}\in A$ as claimed.
Consider now a sequence $\phi_{i}\in A$ minimizing the length functional. Each image of $\phi_{i}$ has a projection on $\gamma_{1}$ lying in $I_{1}^{\gamma_{1}}$. Since $X$ is $(R,D)$-thin, the image of each $\phi_{i}$ has a point which is $D$-close to $I_{1}^{\gamma_{1}}$. Hence, all the images of the maps $\phi_{i}$ are contained in a bounded set of $X$. Since the $\phi_{i}$ are all $1$-Lipschitz, by Ascoli-Arzelà theorem, there exists (modulo passing to a subsequence) $\phi_{\infty}$ such that $\phi_{i}\to\phi_{\infty}$ uniformly. By the previous part of the proof, $\phi_{\infty}\in A$. Since the length functional is lower semicontinuous w.r.t. uniform convergence, $\phi_{\infty}$ is an element of minimal length in $A$.
@@
Step 1. We claim that, for $b\in\{\gamma_{1},\gamma_{2},l\}$, the projection $\pi_{b}(\phi_{\infty}(S^{1}_{L}))$ contains a $3D$-net of $[b(3R),b(\mathrm{L}(b)-3R)]$.
Let $b_{1},b_{2}\in\{\gamma_{1},\gamma_{2},l\}$ with $b_{1}\neq b_{2}$. Let $t^{b_{1}},t^{b_{2}}\in[0,T]$ be the numbers given in the definition of $A$ relative to $\phi_{\infty}$. Modulo reparametrizing $\phi_{\infty}$ and $\gamma_{1}+\gamma_{2}+l$, we can assume that $b_{1}(\mathrm{L}(b_{1}))=b_{2}(0)$ and $t^{b_{1}}\leq t^{b_{2}}$ with $t^{b_{3}}\notin[t^{b_{1}},t^{b_{2}}]$, i.e. $[t^{b_{1}},t^{b_{2}}]$ corresponds to the arc of $S^{1}_{L}$ that connects $t^{b_{1}}$ and $t^{b_{2}}$ without crossing $t^{b_{3}}$.
Let $b_{1}(\bar{t}^{b_{1}})\in I^{b_{1}}$ be a projection on $b_{1}$ of $\phi_{\infty}(t^{b_{1}})$. Similarly, let $b_{2}(\bar{t}^{b_{2}})\in I^{b_{2}}$ be a projection on $b_{2}$ of $\phi_{\infty}(t^{b_{2}})$.
We show that $\pi_{b_{1}}(\phi_{\infty}(S^{1}_{L}))$ contains a $3D$-net of $[b_{1}(\bar{t}^{b_{1}}),b_{1}(\mathrm{L}(b_{1})-3R)]$, while $\pi_{b_{2}}(\phi_{\infty}(S^{1}_{L}))$ contains a $3D$-net of $[b_{2}(3R),b_{2}(\bar{t}^{b_{2}})]$. The claim then follows by repeating the same argument for all pairs of segments in $\{\gamma_{1},\gamma_{2},l\}$.
By definition of $\phi_{\infty}$, the restriction ${\phi_{\infty}}_{|[t^{b_{1}},t^{b_{2}}]}$ is a segment. By Lemmas 4.13 and 4.5, $\pi_{b_{1}}({\phi_{\infty}}([t^{b_{1}},t^{b_{2}}]))$ contains a $3D$-net of either $[b_{1}(3R),b_{1}(\bar{t}^{b_{1}})]$ or $[b_{1}(\bar{t}^{b_{1}}),b_{1}(\mathrm{L}(b_{1})-3R)]$. If the former case happens, ${\phi_{\infty}}([t^{b_{1}},t^{b_{2}}])$ has a point which is $7D$-close to $b_{1}(3R)$, so that it also contains a point $q$ which is $5R$-close to $b_{1}(0)$. By Corollary 4.14, $b_{1}(0)$ is at least $\min\{\mathrm{L}(b_{1}),\mathrm{L}(b_{3})\}+\mathrm{L}(b_{2})/2-20R$ distant from $\phi_{\infty}(t^{b_{2}})$. Hence,
@@
Hence, $\pi_{b_{1}}({\phi_{\infty}}([t^{b_{1}},t^{b_{2}}]))$ contains a $3D$-net of $[b_{1}(\bar{t}^{b_{1}}),b_{1}(\mathrm{L}(b_{1})-3R)]$. The same argument shows that $\pi_{b_{2}}({\phi_{\infty}}([t^{b_{1}},t^{b_{2}}]))$ contains a $3D$-net of $[b_{2}(3R),b_{2}(\bar{t}^{b_{2}})]$. This concludes the proof of Step 1.
@@
As before, w.l.o.g. we can assume that $t^{b_{1}}\leq t^{b_{2}}$. Consider the segment ${\phi_{\infty}}_{|[t^{b_{1}},t^{b_{2}}]}$. We assume by contradiction that there exists a curve $\alpha$ connecting $\phi_{\infty}(t^{b_{1}}+3R)$ and $\phi_{\infty}(t^{b_{2}}+3R)$ in $X$ which does not intersect $B_{2R}(\phi_{\infty}(t^{b_{1}}))\cup B_{2R}(\phi_{\infty}(t^{b_{2}}))$. By item 1 of Lemma 4.5 applied to the segment $\phi_{\infty}([t^{b_{1}},t^{b_{2}}])$ (this can be applied thanks to step $2$), $\pi_{\phi_{\infty}([t^{b_{1}},t^{b_{2}}])}(\alpha)$ contains a $3D$-net of either $[\phi_{\infty}(t^{b_{1}}+R),\phi_{\infty}(t^{b_{1}}+3R)]$ or $[\phi_{\infty}(t^{b_{1}}+3R),\phi_{\infty}(t^{b_{2}}-R)]$. In both cases $\alpha$ intersects $B_{2R}(\phi_{\infty}(t^{b_{1}}))\cup B_{2R}(\phi_{\infty}(t^{b_{2}}))$, a contradiction.
@@
A consequence of Theorem 4.16 is that $(R,D)$-thin metric spaces have $1$-Urysohn width bounded from above by a constant $c(R)>0$ depending only on $R$ (see Corollary 4.18). We recall that the notion of Urysohn width was introduced by Gromov in [FillingGromov, widthGromov, LargeGromov].
@@
$f(x):=\begin{cases}\mathsf{d}(B,x)&\text{if }x\text{ has a projection on }\phi(I)\text{ lying in }\phi(\mathbb{R}_{+})\\ -\mathsf{d}(B,x)&\text{if }x\text{ has a projection on }\phi(I)\text{ lying in}\phi(\mathbb{R}_{-}).\end{cases}$
First of all, $f$ is well defined, since if $x$ has projections on $\phi(I)$ both in ϕ(+)italic-$\phi(\mathbb{R}_{+})$ and in ϕ()italic-$\phi(\mathbb{R}_{-})$, then it must have a projection in $\phi([-200R,200R])$, so that $x\in B$. We now check that $f$ is continuous. This is trivial at points in $\bar{B}$, so let us assume that $x\notin\bar{B}$ and, without loss of generality, that $x$ has a projection on $\phi(I)$ lying in ϕ(+)italic-$\phi(\mathbb{R}_{+})$. Let $x_{i}\to x$. Since $x\notin\bar{B}$, $x$ has a projection on $\phi(I)$ lying in $\phi([500R,+\infty))$ so that, for $i$ large enough, $x_{i}$ has a projection on $\phi(I)$ lying in ϕ(+)italic-$\phi(\mathbb{R}_{+})$. It follows that $f$ is continuous at $x$, which implies that $f$ is continuous.
Let now $t\in\mathbb{R}$ and let $x,y\in f^{-1}(t)$. If $t=0$, then $\mathsf{d}(x,y)\leq 1000R$. Assume now that $t>0$, and let $t_{x},t_{y}\in\mathbb{R}$ be such that $\phi(t_{x})$ and $\phi(t_{y})$ are projections of $x$ and $y$ on $\phi(I)$. Since $t>0$, it follows that $t_{x},t_{y}>0$, so that $t_{x},t_{y}\in[\mathsf{d}(\phi(0),x)-200R,\mathsf{d}(\phi(0),x)+200R]$, which implies $\mathsf{d}(x,y)\leq 1000R$. The case $t<0$ is analogous, concluding the case $I=\mathbb{R}$.
Case 3: $I=S^{1}_{L}=[0,2\pi L]/\sim$, where $\sim$ identifies the endpoints of $[0,2\pi L]$.
If $L\leq 10^{5}R$, the zero map $f:X\to\mathbb{R}$ defined as $f(x):=0$ gives the desired $1$-Urysohn width estimate. So, we can assume $L\geq 10^{5}R$. Let $B_{1}:=B_{1000R}(\phi(0))$, $B_{2}:=B_{1000R}(\phi(\pi L))$, and consider the map $f:X\to[0,2\mathsf{d}(B_{1},B_{2})]/\sim$, where $\sim$ identifies the endpoints of $[0,2\mathsf{d}(B_{1},B_{2})]$, given by
$f(x):=\begin{cases}\mathsf{d}(B_{1},x)\wedge\mathsf{d}(B_{1},B_{2})&\text{if }x\text{ has a projection on }\phi(I)\text{ lying in }\phi([0,\pi L])\\ (\mathsf{d}(B_{1},B_{2})+\mathsf{d}(B_{2},x))\wedge 2\mathsf{d}(B_{1},B_{2})&\text{if }x\text{ has a projection on }\phi(I)\text{ lying in }\\ &\;\phi([\pi L,2\pi L]).\end{cases}$
The map $f$ is well defined, since if $x$ has projections on $\phi(I)$ both in $\phi([0,\pi L])$ and in $\phi([\pi L,2\pi L])$, then $x\in B_{1}\cup B_{2}$. We now check that $f$ is continuous. This is trivial at points in $B_{1}\cup B_{2}$. So, we consider $x\notin B_{1}\cup B_{2}$ and we assume, without loss of generality, that $x$ has a projection on $\phi(I)$ lying in $\phi([0,\pi L])$. Let $x_{i}\to x$. Since $x\notin B_{1}\cup B_{2}$, $x$ has a projection on $\phi(I)$ lying in $\phi([500R,\pi L-500R])$, so that, for $i$ large enough, $x_{i}$ has a projection on $\phi(I)$ lying in $\phi([0,\pi L])$. It follows that $f$ is continuous at $x$, so that $f$ is continuous.
@@ 5.1 Sufficient curvature conditions for a manifold to be thin
The next theorem studies spaces of the form $\mathbb{R}^{n-3}\times X$ arising as limits of $n$-manifolds with almost non-negative Ricci curvature and positive scalar curvature in integral sense. One should compare also this result with [wang2023positive, Theorem 1.1], where limit spaces of the form $\mathbb{R}^{n-2}\times X$ were studied.
@@
In addition, $X$ contains a line which arises as a limit of the segments $r_{j}$ centered in $r_{j}(R_{j})$, since $R_{j}\uparrow\infty$. Hence, by the splitting theorem [Gigli-Split], $X$ splits isomorphically as a metric measure space as $X=Y\times\mathbb{R}$ for an $\mathsf{RCD}(0,n-1)$ space $(Y,\mathsf{d}_{y},\mathfrak{m}_{y})$. Moreover, $Y$ cannot be compact since the segments $\gamma_{j}$, which realize the distance from $r_{j}$, have arbitrarily large length. In particular, $Y$ contains itself a ray. Hence, repeating the argument above, there exists a sequence $p_{j}^{\prime}\in Y$, such that $(Y,\mathsf{d}_{y},p_{j}^{\prime})$ converges in pGH-sense to a metric space ×$\mathbb{R}\times Y^{\prime}$. In particular, there exists a sequence $p_{j}^{\prime\prime}\in X_{j}$ such that $(X_{j},\mathsf{d}_{j},p_{j}^{\prime\prime})$ converges in pGH-sense to $\mathbb{R}^{2}\times Y^{\prime}$. Applying Theorem 1.1 to $M_{i,j}$ for $i$ and $j$ large enough, we arrive to a contradiction, since $\delta_{j}\downarrow 0$ and $\epsilon_{0}>0$ is fixed in (37).
@@
The remaining part of the section is devoted to showing that if a manifold satisfies one of the conditions (34), (35), (36), so do its coverings (modulo a rescaling by dimensional constants). This is the content of Propositions 5.3 and 5.4. The proof of the next proposition follows from [kapovitch2011structurefundamentalgroupsmanifolds, Lemma 1.6]; we report it for completeness of presentation.
@@
Let $n\in\mathbb{N}$ and let $K\in\mathbb{R}$. Let $(M^{n},g,p)$ be a pointed Riemannian manifold with $\mathsf{Ric}_{M}\geq K$. Let $(\tilde{M}^{n},\tilde{g},\tilde{p})$ be the universal covering of $M$, with covering map $\pi:\tilde{M}\to M$ such that $\pi(\tilde{p})=p$. Let $F:M\to\mathbb{R}_{+}$ be a measurable locally bounded function and let $\tilde{F}:=F\circ\pi$.
@@
Consider a measurable function $j:B_{1}(p)\to\tilde{B}_{1}(\tilde{p})$ such that $\pi(j(x))=x$ for every $x\in B_{1}(p)$. Let $T:=j(B_{1}(p))$, so that $\mathrm{diam}(T)\leq 2$ and
@@
Let $S\subset\tilde{M}$ be the union of the sets $g(T)$ for every deck transformation $g:\tilde{M}\to\tilde{M}$ of $\pi:\tilde{M}\to M$ such that $g(T)\cap\tilde{B}_{1}(\tilde{p})\neq\emptyset$. It follows that
@@
Proof of 1. Using that $\widetilde{\mathsf{Vol}}(\tilde{B}_{3}(\tilde{p}))\leq c(K,n)\widetilde{\mathsf{Vol}}(\tilde{B}_{1}(\tilde{p}))$ and (41), it holds
[…]
@@
[…]
@@
[…]
By Bishop-Gromov’s inequality, it holds #$\#I\leq c(K,n)$. Setting $p^{\prime}\in\{p_{i}\}_{i\in I}$ such that
[…]
@@
Let $v,\epsilon,\delta,L\in(0,+\infty)$, $n\in\mathbb{N}$, $s\in(0,1)$ be fixed. Let $(M^{n},g,p)$ be a manifold satisfying one of the conditions (34), (35), (36). Let $(\tilde{M}^{n},\tilde{g},\tilde{p})$ be the universal covering, with covering map $\pi:\tilde{M}\to M$ such that $\pi(\tilde{p})=p$. Then $(\tilde{M},\tilde{g}/3^{2},\tilde{p})$ satisfies the corresponding condition among (34), (35), (36), where the constants have been rescaled by factors depending only on the dimension $n$.
Corollary 5.5.
Let $v,\epsilon,L\in(0,+\infty)$, $n\in\mathbb{N}$, $s\in(0,1)$ be fixed. There exist $R,D,\delta>0$ with $R\geq 20D$ satisfying the following. Let $(M^{n},g)$ be a manifold satisfying one of the conditions (38), (39), (40). Let $(\tilde{M},\tilde{g},\tilde{p})$ be the universal covering of $M$ and let $\pi:\tilde{M}\to M$ be the covering map. Then, both $(M,g)$ and $(\tilde{M},\tilde{g})$ are $(R,D)$-thin (see Definition 4.2).

### v3

@@ 4 Thin metric spaces
$\quad\pi_{b}\big(\phi(t^{b})\big)\cap I^{b}\neq\emptyset\text{ for every }b\in\{\gamma_{1},\gamma_{2},l\}\Big\}.$
Notice that $A\neq\emptyset$ since $\gamma_{1}+\gamma_{2}+l\in A$. We claim that if $\{\phi_{i}\}_{i\in\mathbb{N}}\subset A$ and $\phi_{i}\to\phi_{\infty}$ uniformly, then $\phi_{\infty}\in A$. First of all $\phi_{\infty}$ is $1$-Lipschitz and satisfies $\phi_{\infty}(0)=\phi_{\infty}(\gamma_{1}+\gamma_{2}+l)$ trivially. Let $b\in\{\gamma_{1},\gamma_{2},l\}$. Consider for every $i\in\mathbb{N}$ the corresponding $t_{i}^{b}$. Modulo passing to a subsequence, there exists $t_{\infty}^{b}\in[0,\mathrm{L}(\gamma_{1}+\gamma_{2}+l)]$ such that $t_{i}^{b}\to t_{\infty}^{b}$. By Lemma 4.4, $\pi_{b}\big(\phi_{\infty}(t_{\infty}^{b})\big)\cap I^{b}\neq\emptyset$. It follows that $\phi_{\infty}\in A$ as claimed.
Consider now a sequence $\phi_{i}\in A$ minimizing the length functional. Each image of $\phi_{i}$ satisfies $\pi_{\gamma_{1}}\big(\mathrm{Im}(\phi_{i})\big)\cap I_{1}^{\gamma_{1}}\neq\emptyset$. Since $X$ is $(R,D)$-thin, the image of each $\phi_{i}$ has a point which is $D$-close to $I_{1}^{\gamma_{1}}$. Hence, all the images of the maps $\phi_{i}$ are contained in a bounded set of $X$. Since the $\phi_{i}$ are all $1$-Lipschitz, by Ascoli-Arzelà theorem, there exists (modulo passing to a subsequence) $\phi_{\infty}$ such that $\phi_{i}\to\phi_{\infty}$ uniformly. By the previous part of the proof, $\phi_{\infty}\in A$. Since the length functional is lower semicontinuous w.r.t. uniform convergence, $\phi_{\infty}$ is an element of minimal length in $A$.
@@
Step 1. We claim that, for $b\in\{\gamma_{1},\gamma_{2},l\}$, the projection $\pi_{b}(\phi_{\infty}(S^{1}_{L}))$ contains a $3D$-net of $b(3R),b(\mathrm{L}(b)-3R)]$.
Let $b_{1},b_{2}\in\{\gamma_{1},\gamma_{2},l\}$ with $b_{1}\neq b_{2}$. Let $t^{b_{1}},t^{b_{2}}\in[0,T]$ be the numbers given in the definition of $A$ relative to $\phi_{\infty}$. Modulo reparametrizing $\phi_{\infty}$ and $\gamma_{1}+\gamma_{2}+l$, we can assume that $b_{1}(\mathrm{L}(b_{1}))=b_{2}(0)$ and $t^{b_{1}}\leq t^{b_{2}}$ with $t^{b_{3}}\notin[t^{b_{1}},t^{b_{2}}]$, i.e. $t^{b_{1}},t^{b_{2}}]$ corresponds to the arc of $S^{1}_{L}$ that connects $t^{b_{1}}$ and $t^{b_{2}}$ without crossing $t^{b_{3}}$.
Let $b_{1}(\bar{t}^{b_{1}})\in I^{b_{1}}\cap\pi_{b_{1}}\big(\phi_{\infty}(t^{b_{1}})\big)$. Similarly, let $b_{2}(\bar{t}^{b_{2}})\in I^{b_{2}}\cap\pi_{b_{2}}\big(\phi_{\infty}(t^{b_{2}})\big)$. We show that $\pi_{b_{1}}(\phi_{\infty}(S^{1}_{L}))$ contains a $3D$-net of $b_{1}(\bar{t}^{b_{1}}),b_{1}(\mathrm{L}(b_{1})-3R)]$, while $\pi_{b_{2}}(\phi_{\infty}(S^{1}_{L}))$ contains a $3D$-net of $b_{2}(3R),b_{2}(\bar{t}^{b_{2}})]$. The claim then follows by repeating the same argument for all pairs of segments in $\{\gamma_{1},\gamma_{2},l\}$.
By definition of $\phi_{\infty}$, the restriction ${\phi_{\infty}}_{|[t^{b_{1}},t^{b_{2}}]}$ is a segment. By Lemmas 4.13 and 4.5, $\pi_{b_{1}}({\phi_{\infty}}([t^{b_{1}},t^{b_{2}}]))$ contains a $3D$-net of either $b_{1}(3R),b_{1}(\bar{t}^{b_{1}})]$ or $b_{1}(\bar{t}^{b_{1}}),b_{1}(\mathrm{L}(b_{1})-3R)]$. If the former case happens, ${\phi_{\infty}}([t^{b_{1}},t^{b_{2}}])$ has a point which is $7D$-close to $b_{1}(3R)$, so that it also contains a point $q$ which is $5R$-close to $b_{1}(0)$. By Corollary 4.14, $b_{1}(0)$ is at least $\min\{\mathrm{L}(b_{1}),\mathrm{L}(b_{3})\}+\mathrm{L}(b_{2})/2-20R$ distant from $\phi_{\infty}(t^{b_{2}})$. Hence,
@@
Hence, $\pi_{b_{1}}({\phi_{\infty}}([t^{b_{1}},t^{b_{2}}]))$ contains a $3D$-net of $b_{1}(\bar{t}^{b_{1}}),b_{1}(\mathrm{L}(b_{1})-3R)]$. The same argument shows that $\pi_{b_{2}}({\phi_{\infty}}([t^{b_{1}},t^{b_{2}}]))$ contains a $3D$-net of $b_{2}(3R),b_{2}(\bar{t}^{b_{2}})]$. This concludes the proof of Step 1.
@@
As before, w.l.o.g. we can assume that $t^{b_{1}}\leq t^{b_{2}}$. Consider the segment ${\phi_{\infty}}_{|[t^{b_{1}},t^{b_{2}}]}$. We assume by contradiction that there exists a curve $\alpha$ connecting $\phi_{\infty}(t^{b_{1}}+3R)$ and $\phi_{\infty}(t^{b_{2}}+3R)$ in $X$ which does not intersect $B_{2R}(\phi_{\infty}(t^{b_{1}}))\cup B_{2R}(\phi_{\infty}(t^{b_{2}}))$. By item 1 of Lemma 4.5 applied to the segment $\phi_{\infty}([t^{b_{1}},t^{b_{2}}])$ (this can be applied thanks to step $2$), $\pi_{\phi_{\infty}([t^{b_{1}},t^{b_{2}}])}(\alpha)$ contains a $3D$-net of either $\phi_{\infty}(t^{b_{1}}+R),\phi_{\infty}(t^{b_{1}}+3R)]$ or $\phi_{\infty}(t^{b_{1}}+3R),\phi_{\infty}(t^{b_{2}}-R)]$. In both cases $\alpha$ intersects $B_{2R}(\phi_{\infty}(t^{b_{1}}))\cup B_{2R}(\phi_{\infty}(t^{b_{2}}))$, a contradiction.
@@
A consequence of Theorem 4.16 is that $(R,D)$-thin metric spaces have $1$-Urysohn width bounded from above by a constant $c(R)>0$ depending only on $R$ (see Corollary 4.18). We recall that the notion of Urysohn width was introduced by Gromov in [30, 27, 31].
@@
$f(x):=\begin{cases}\mathsf{d}(B,x)&\text{if }\pi_{\phi(I)}(x)\cap\phi(\mathbb{R}_{+})\neq\emptyset\\ -\mathsf{d}(B,x)&\text{if }\pi_{\phi(I)}(x)\cap\phi(\mathbb{R}_{-})\neq\emptyset.\end{cases}$
First of all, $f$ is well-defined, since if $\pi_{\phi(I)}(x)\cap\phi(\mathbb{R}_{+})\cap\phi(\mathbb{R}_{-})\neq\emptyset$, then
$\pi_{\phi(I)}(x)\cap\phi([-200R,200R])\neq\emptyset,$
so that $x\in B$. We now check that $f$ is continuous. This is trivial at points in $\bar{B}$, so let us assume that $x\notin\bar{B}$ and, without loss of generality, that $\pi_{\phi(I)}(x)\cap\phi(\mathbb{R}_{+})\neq\emptyset$. Let $x_{i}\to x$. Since $x\notin\bar{B}$, $\pi_{\phi(I)}(x)\cap\phi([500R,+\infty))\neq\emptyset$ so that, for $i$ large enough, $\pi_{\phi(I)}(x_{i})\cap\phi(\mathbb{R}_{+})\neq\emptyset$. It follows that $f$ is continuous at $x$, which implies that $f$ is continuous.
Let now $t\in\mathbb{R}$ and let $x,y\in f^{-1}(t)$. If $t=0$, then $\mathsf{d}(x,y)\leq 1000R$. Assume now that $t>0$, and let $t_{x},t_{y}\in\mathbb{R}$ be such that $\phi(t_{x})\in\pi_{\phi(I)}(x)$ and $\phi(t_{y})\in\pi_{\phi(I)}(y)$. Since $t>0$, it follows that $t_{x},t_{y}>0$, so that $t_{x},t_{y}\in[\mathsf{d}(\phi(0),x)-200R,\mathsf{d}(\phi(0),x)+200R]$, which implies $\mathsf{d}(x,y)\leq 1000R$. The case $t<0$ is analogous, concluding the case $I=\mathbb{R}$.
Case 3: $I=S^{1}_{L}=[0,2\pi L]/\sim$, where $\sim$ identifies the endpoints of $0,2\pi L]$.
If $L\leq 10^{5}R$, the zero map $f:X\to\mathbb{R}$ defined as $f(x):=0$ gives the desired $1$-Urysohn width estimate. So, we can assume $L\geq 10^{5}R$. Let $B_{1}:=B_{1000R}(\phi(0))$, $B_{2}:=B_{1000R}(\phi(\pi L))$, and consider the map $f:X\to[0,2\mathsf{d}(B_{1},B_{2})]/\sim$, where $\sim$ identifies the endpoints of $0,2\mathsf{d}(B_{1},B_{2})]$, given by
$f(x):=\begin{cases}\mathsf{d}(B_{1},x)\wedge\mathsf{d}(B_{1},B_{2})&\text{if }\pi_{\phi(I)}(x)\cap\phi([0,\pi L])\neq\emptyset\\ (\mathsf{d}(B_{1},B_{2})+\mathsf{d}(B_{2},x))\wedge 2\mathsf{d}(B_{1},B_{2})&\text{if }\pi_{\phi(I)}(x)\cap\phi([\pi L,2\pi L])\neq\emptyset.\end{cases}$
The map $f$ is well-defined, since if $\pi_{\phi(I)}(x)\cap\phi([0,\pi L])\cap\phi([\pi L,2\pi L])\neq\emptyset$, then $x\in B_{1}\cup B_{2}$. We now check that $f$ is continuous. This is trivial at points in $B_{1}\cup B_{2}$. So, we consider $x\notin B_{1}\cup B_{2}$ and we assume, without loss of generality, that $\pi_{\phi(I)}(x)\cap\phi([0,\pi L])\neq\emptyset$. Let $x_{i}\to x$. Since $x\notin B_{1}\cup B_{2}$, $\pi_{\phi(I)}(x)\cap\phi([500R,\pi L-500R])\neq\emptyset$, so that, for $i$ large enough, $\pi_{\phi(I)}(x_{i})\cap\phi([0,\pi L])\neq\emptyset$. It follows that $f$ is continuous at $x$, so that $f$ is continuous.
@@ 5.1 Sufficient curvature conditions for a manifold to be thin
The next theorem studies spaces of the form $\mathbb{R}^{n-3}\times X$ arising as limits of $n$-manifolds with almost non-negative Ricci curvature and positive scalar curvature in integral sense. One should compare also this result with [59, Theorem 1.1], where limit spaces of the form $\mathbb{R}^{n-2}\times X$ were studied.
@@
In addition, $X$ contains a line which arises as a limit of the segments $r_{j}$ centered in $r_{j}(R_{j})$, since $R_{j}\uparrow\infty$. Hence, by the splitting theorem [24], $X$ splits isomorphically as a metric measure space as $X=Y\times\mathbb{R}$ for an $\mathsf{RCD}(0,n-1)$ space $(Y,\mathsf{d}_{y},\mathfrak{m}_{y})$. Moreover, $Y$ cannot be compact since the segments $\gamma_{j}$, which realize the distance from $r_{j}$, have arbitrarily large length. In particular, $Y$ contains itself a ray. Hence, repeating the argument above, there exists a sequence $p_{j}^{\prime}\in Y$, such that $(Y,\mathsf{d}_{y},p_{j}^{\prime})$ converges in pGH-sense to a metric space $\mathbb{R}\times Y^{\prime}$. In particular, there exists a sequence $p_{j}^{\prime\prime}\in X_{j}$ such that $(X_{j},\mathsf{d}_{j},p_{j}^{\prime\prime})$ converges in pGH-sense to $\mathbb{R}^{2}\times Y^{\prime}$. Applying Theorem 1.1 to $M_{i,j}$ for $i$ and $j$ large enough, we arrive to a contradiction, since $\delta_{j}\downarrow 0$ and $\epsilon_{0}>0$ is fixed in (29).
@@
The remaining part of the section is devoted to showing that if a manifold satisfies one of the conditions (26), (27), (28), so do its coverings (modulo a rescaling by dimensional constants). This is the content of Propositions 5.3 and 5.4. The proof of the next proposition follows from [35, Lemma 1.6]; we report it for completeness of presentation.
@@
[…]
@@
$\fint_{T}\tilde{F}\,d\widetilde{\mathsf{Vol}}=\fint_{B_{1}(p)}F\,d\mathsf{Vol}.$
@@
$\tilde{B}_{1}(\tilde{p})\subset S\subset\tilde{B}_{3}(\tilde{p})\quad\text{and}\quad\fint_{S}\tilde{F}\,d\widetilde{\mathsf{Vol}}=\fint_{B_{1}(p)}F\,d\mathsf{Vol}.$ (33)
Proof of 1. Using that $)\widetilde{\mathsf{Vol}}(\tilde{B}_{3}(\tilde{p}))\leq c(K,n)\widetilde{\mathsf{Vol}}(\tilde{B}_{1}(\tilde{p}$)) and (33), it holds
$\fint_{B_{1}(p)}F\,d\mathsf{Vol}\leq\frac{\widetilde{\mathsf{Vol}}(\tilde{B}_{3}(\tilde{p}))}{\widetilde{\mathsf{Vol}}(\tilde{B}_{1}(\tilde{p}))}\fint_{\tilde{B}_{3}(\tilde{p})}\tilde{F}\,d\widetilde{\mathsf{Vol}}\leq c(K,n)\fint_{\tilde{B}_{3}(\tilde{p})}\tilde{F}\,d\widetilde{\mathsf{Vol}}$.
@@
$\tilde{B}_{3}(\tilde{p})\subset S^{\prime}\subset\tilde{B}_{9}(\tilde{p})\quad\text{and}\quad\fint_{S^{\prime}}\tilde{F}\,d\widetilde{\mathsf{Vol}}=\fint_{B_{3}(p)}F\,d\mathsf{Vol}.$ (34)
@@
$\fint_{\tilde{B}_{3}(\tilde{p})}\tilde{F}\,d\widetilde{\mathsf{Vol}}\leq\frac{\widetilde{\mathsf{Vol}}(\tilde{B}_{9}(\tilde{p}))}{\widetilde{\mathsf{Vol}}(\tilde{B}_{3}(\tilde{p}))}\fint_{B_{3}(p)}F\,d\mathsf{Vol}\leq c(K,n)\sum_{i\in I}\fint_{B_{1}(p_{i})}F\,d\mathsf{Vol}.$ (35)
By Bishop-Gromov’s inequality, it holds $I\leq c(K,n)$. Setting $p^{\prime}\in\{p_{i}\}_{i\in I}$ such that
$\fint_{B_{1}(p^{\prime})}F\,d\mathsf{Vol}=\max_{i\in I}\fint_{B_{1}(p_{i})}F\,d\mathsf{Vol},$
@@
[…]
Corollary $5.$
[…]